the sum of an even integer and twice it's consecutive even integeris 76.what are the integers? please translate this into consecutive number problems

Answer by mananth(16949)

let the even integer be x
the next will be x+2
x+2(x+2)=76
x+2x+4=76
3x=76-4
3x=72
x=24
The numbers are 24 and 26